# ExaGRyPE: Numerical General Relativity Solvers Based upon the Hyperbolic PDEs Solver Engine ExaHyPE

### v2

@@ 5.3 Rotating Binary Black Hole Merger
$\alpha=\frac{1}{2}\left(\frac{1-\frac{1}{2}\left(m_{-}/r_{-}\right)-\frac{1}{2}\left(m_{+}/r_{+}\right)}{1+\frac{1}{2}\left(m_{-}/r_{-}\right)+\frac{1}{2}\left(m_{+}/r_{+}\right)}+1\right),$
with vanished initial shift is used. Otherwise, we preserve the black hole settings, i.e. $\kappa_{1}=0.1,\kappa_{2}=0,\kappa_{3}=0.5,e=c=\tau=1.0$ and $\mu=0.2$. The gamma driver gauge condition uses the parameters $f=0.75$ and $\eta=1$. Compared to the FD4 setup, no modifications are made to the KO coefficient ($8$) and the CFL ratio ($0.1$).
We employ a domain of $[-12M,12M]^{3}$ with three levels of AMR. The domain gets refined at radius $r=9M$ and again at $r=5M$ around the center of the coordinate system, where we assume the black holes to merge. We obtain FD spacings at each level of $[0.333M,0.111M,0.037M]$. Sommerfeld boundary conditions are used. We run both the first-order and second-order formulation for this test, taking into account previously described stability challenges for the former once the black holes have merged.
@@
Figure 16: The trajectories of the two punctures in the binary black hole merger test. The two black holes $\pm$ are initially located at $[x,y,z]=[2M,0,0],[-2M,0,0]$.
The binary black holes complete one-and-a-half circular orbits before they merge at a code time of $T\approx 70$. The system continues to evolve as a single black hole until $T\approx 120$, where we terminate the simulation. The rotation and merging processes in the simulation are relatively smooth and subject to only limited violations of the constraints (Figures 14 and 15). As the punctures are rotating, the shift vector exhibits a spiral pattern. Its reversed direction at the puncture location indicates the velocity direction of the puncture, i.e., we can use it to inform ExaGRyPE’s tracers from Section 3.8 (Figure 16). The spiral pattern in the post-merge stage shows that the remnant of the merged black holes carries a spin. Besides the actual physical data, some fluctuation starts to creep into the simulation. It stems from the resolution transitions and gets magnified in the post-merge phase. In line with previous experiments, the first-order formulation encounters instability after merging and soon runs into a crash. However, it can properly resolve the rotation of black holes before that. The dynamic behavior of the first-order and second-order scheme are presented in Figure 17, both schemes have the same trajectories until the approaching punctures accumulate errors and start to affect the evolution.
Overall, the rotation and pre-merger phase are accurately captured, and the locations of the punctures are correctly tracked by the (inverse) shift vector guiding the tracers. The simulation of the dynamic behavior is stable, which supports the argument that any constraint violation is “traveling” through the domain, might be reflected at AMR boundaries, but eventually leaves the domain. We avoid the accumulation of numerical errors at specific locations as long as the black holes move.
@@ 5.4 Context of Experimental Studies
Our application successfully completes a rotating binary black hole merger. It is able to handle both the approach phase as well as the actual ringdown. Our studies focus on the introduction of numerical ingredients within ExaGRyPE. They validate that the software is a release candidate, ready for larger production runs. Here, we have to combine all ingredients and notably study a larger computational domain to avoid boundary artifacts (cmp. squeeze-in effect in Figure 17) and obtain realistic measurements. Further stationery tracers (cmp. Section 5.1) then can be used to sample the solution on a huge sphere around the area of interest. With such samples, we can finally calculate $\psi_{4}$ and its mode quantities.
@@ 6 Discussion and Conclusion
To implement this vision and to showcase the applicability and maturity of the design, the ExaGRyPE endeavor had to add further building blocks to the underlying ExaHyPE engine, and we had to run an extensive set of numerical experiments. The experiments demonstrate that our code, in principle, is ready for production, but also raise further questions, notably around the numerical stability of any composition of ingredients and the philosophical question if first- or second-order formulations are superior. Many aspects of these questions are intrinsically tied to this area of research and have been documented, either implicitly or explicitly, by other codes and groups, too. We consider it a strength of ExaGRyPE that we can quite flexibly study and swap different methodological ingredients to address these fundamental research questions, although the present work does not yet provide answers to many of them.
@@ Appendix B Spacetime Foliation and Evolution Equations
One can already conduct simulations in this stage as $\gamma_{ij},K_{ij},\alpha,\beta^{i}$ with $Z^{a}$ field form a complete evolution system. However, it is proven to be unstable with a low hyperbolicity. Therefore, a conformal decomposition is applied to the system [11], where we define a conformal factor $\phi$ and a conformal metric $\tilde{\gamma}_{ij}$, such that
@@
$\begin{split}&\partial_{t}\hat{\Gamma}^{i}-\beta^{k}\partial_{k}\hat{\Gamma}^{i}-2\alpha\tilde{\gamma}^{ki}\partial_{k}\Theta-\tilde{\gamma}^{kl}\partial_{(k}B_{l)}^{i}-s\frac{1}{3}\tilde{\gamma}^{ik}\partial_{(k}B_{l)}^{l}\\ &\quad\quad-{\color[rgb]{1,0,0}2\alpha\tilde{\gamma}^{ik}\tilde{\gamma}^{nm}\partial_{k}\tilde{A}_{nm}}+\frac{4}{3}\alpha\tilde{\gamma}^{ij}\partial_{j}K=\frac{2}{3}\tilde{\Gamma}^{i}B_{k}^{k}-\tilde{\Gamma}^{k}B_{k}^{i}\\ &\quad\quad+2\alpha\left(\tilde{\Gamma}_{jk}^{i}\tilde{A}^{jk}-3\tilde{A}^{ij}\frac{P_{j}}{\phi}\right)-2\alpha\tilde{\gamma}^{ki}\left(\Theta\frac{A_{k}}{\alpha}+\frac{2}{3}KZ_{k}\right)\\ &\quad\quad-2\tilde{A}^{ij}A_{j}-{\color[rgb]{1,0,0}4\alpha\tilde{\gamma}^{ik}D_{k}^{nm}\tilde{A}_{nm}}-2\alpha\kappa_{1}\tilde{\gamma}^{ij}Z_{j}\\ &\quad\quad+2\kappa_{3}\left(\frac{2}{3}\tilde{\gamma}^{ij}Z_{j}B_{k}^{k}-\tilde{\gamma}^{jk}Z_{j}B_{k}^{i}\right),\end{split}$ (28)
@@
$\begin{split}&\partial_{t}A_{k}-\beta^{l}\partial_{l}A_{k}+\alpha^{2}g(\alpha)\left(\partial_{k}K-\partial_{k}K_{0}-2c\partial_{k}\Theta\right)\\ &\qquad+{\color[rgb]{1,0,0}\alpha g(\alpha)\tilde{\gamma}^{nm}\partial_{k}\tilde{A}_{nm}}=B^{l}_{k}A_{l}+{\color[rgb]{1,0,0}2\alpha g(\alpha)D_{k}^{\ nm}\tilde{A}_{nm}}\\ &\quad\quad\quad\quad\quad\quad-\left[2{\alpha}g(\alpha)+\alpha^{2}g^{\prime}(\alpha)\right]\left(K-K_{0}-2c\Theta\right)A_{k},\end{split}$ (30)
$\begin{split}&\partial_{t}B_{k}^{i}-\beta^{l}\partial_{l}B_{k}^{i}-f\partial_{k}b^{i}+{\color[rgb]{1,0,0}\alpha^{2}\mu\frac{\tilde{\gamma}^{ij}}{\phi}\left(\partial_{k}P_{j}-\partial_{j}P_{k}\right)}\\ &\quad\quad\quad\quad\quad-{\color[rgb]{1,0,0}\alpha^{2}\mu\tilde{\gamma}^{ij}\tilde{\gamma}^{nl}\left(\partial_{k}D_{ljn}-\partial_{l}D_{kjn}\right)}=B_{k}^{l}B_{l}^{i},\end{split}$ (31)
$\begin{split}&\partial_{t}D_{kij}-\beta^{l}\partial_{l}D_{kij}+s\Big{(}-\frac{1}{2}\tilde{\gamma}_{mi}\partial_{(k}B_{j)}^{m}-\frac{1}{2}\tilde{\gamma}_{mj}\partial_{(k}B_{i)}^{m}\\ &\quad+\frac{1}{3}\tilde{\gamma}_{ij}\partial_{(k}B_{m)}^{m}\Big{)}+\alpha\partial_{k}\tilde{A}_{ij}={\color[rgb]{1,0,0}\alpha\frac{1}{3}\tilde{\gamma}_{ij}\tilde{\gamma}^{nm}\partial_{k}\tilde{A}_{nm}}\\ &\quad+B_{k}^{l}D_{lij}+B_{j}^{l}D_{kli}+B_{i}^{l}D_{klj}-\frac{2}{3}B_{l}^{l}D_{kij}\\ &\quad-{\color[rgb]{1,0,0}\alpha\frac{2}{3}\tilde{\gamma}_{ij}D_{k}^{nm}\tilde{A}_{nm}}-A_{k}\left(\tilde{A}_{ij}-\frac{1}{3}\tilde{\gamma}_{ij}\operatorname{tr}\tilde{A}\right),\end{split}$ (32)
$\begin{split}&\partial_{t}P_{k}-\beta^{l}\partial_{l}P_{k}-\frac{1}{3}\phi\left(\alpha\partial_{k}K-\partial_{(k}B_{l)}^{l}\right)-{\color[rgb]{1,0,0}\frac{1}{3}\alpha\phi\tilde{\gamma}^{nm}\partial_{k}\tilde{A}_{nm}}\\ &\quad=B^{l}_{k}P_{l}+\frac{1}{3}\left(\alpha K-B^{l}_{l}\right)P_{k}+\frac{1}{3}\phi KA_{k}-{\color[rgb]{1,0,0}\frac{2}{3}\alpha\phi D_{k}^{\ nm}\tilde{A}_{nm}}.\end{split}$ (33)
Most of the symbols already appear in the context above and are self-explaining. The $\mathrm{TF}$ index means the Trace-Free part of the quantities, i.e., $R^{\mathrm{TF}}_{ij}=R_{ij}-\frac{1}{3}\gamma_{ij}\gamma^{kl}R_{kl}$. $\operatorname{tr}\tilde{A}=\gamma^{ij}\tilde{A}_{ij}$ is the trace of the conformal traceless extrinsic curvature. There are several red terms added to the evolution systems, using the constraints (20) above for a more symmetric characteristic matrix [10]. Those terms are not necessary in the second-order formulation. The $g(\alpha)$ function is the one defined in the slicing condition in Section 2.1, and the “1+log” slicing corresponds to $g(\alpha)=2/\alpha$. Several new parameters are also introduced for numerical optimization:
@@ Appendix C $\psi_{4}$ in ExaGRyPE
where ${}^{(4)}C_{abcd}$ is the four-dimensional Weyl tensor, and $k^{a}$ and $\bar{m}^{a}$ are two members of a null tetrad. A tetrad is a set of four null vectors that form a complete basic, and we name them $l^{a}$, $k^{a}$, $m^{a}$ and $\bar{m}^{a}$. The tetrad is constructed such that the $l^{a}$ and $k^{a}$ are radial outgoing and ingoing vectors respectively, and only non-vanishing inner products between them are:

### v3

@@ 5.3 Rotating Binary Black Hole Merger
$\alpha=\frac{1}{2}\left(\frac{1-\frac{1}{2}\left(m_{-}/r_{-}\right)-\frac{1}{2}\left(m_{+}/r_{+}\right)}{1+\frac{1}{2}\left(m_{-}/r_{-}\right)+\frac{1}{2}\left(m_{+}/r_{+}\right)}+1\right),$ (39)
with vanished initial shift is used. Otherwise, we preserve the black hole settings, i.e. $\kappa_{1}=0.1,\kappa_{2}=0,\kappa_{3}=0.5,e=c=\tau=1.0$ and $\mu=0.2$. The gamma driver gauge condition uses the parameters $f=0.75$ and $\eta=1$. Compared to the FD4 setup, no modifications are made to the KO coefficient ($8$) and the CFL ratio ($0.1$). We also use the RK1 temporal scheme for this scenario, and convergence tests have confirmed that we achieve the expected order ($\sim h^{1}$) of convergence across different resolutions with it (test not shown).
We employ a domain of $[-12M,12M]^{3}$ with three levels of SMR. The domain gets refined at radius $r=9M$ and again at $r=5M$ around the center of the coordinate system, where we assume the black holes to merge. We obtain FD spacings at each level of $[0.333M,0.111M,0.037M]$. Sommerfeld boundary conditions are used. We run both the first-order and second-order formulation for this test, taking into account previously described stability challenges for the former once the black holes have merged.
@@
Figure 18: The trajectories of the two punctures in the binary black hole merger test. The two black holes $\pm$ are initially located at $[x,y,z]=[2M,0,0],[-2M,0,0]$. The trajectories for the final ring-down stage were removed for clarity. The kink observed in the initial stage of the $+$ puncture arises from a minor inconsistency in data output as the tracer transitions between different subpartitions managed by separate MP threads. The symmetric trajectories that follow confirm that there is no underlying computational error.
The binary black holes complete one-and-a-half circular orbits before they merge at a code time of $T\approx 70$. The system continues to evolve as a single black hole until $T\approx 120$, where we terminate the simulation. The rotation and merging processes in the simulation are relatively smooth and subject to only limited violations of the constraints (Figures 16 and 17). As the punctures are rotating, the shift vector exhibits a spiral pattern. Its reversed direction at the puncture location indicates the velocity direction of the puncture, i.e., we can use it to inform ExaGRyPE’s tracers from Section 3.8 (Figure 18). The spiral pattern in the post-merge stage shows that the remnant of the merged black holes carries a spin. Besides the actual physical data, some fluctuation starts to creep into the simulation. It stems from the resolution transitions and gets magnified in the post-merge phase. In line with previous experiments, the first-order formulation encounters instability after merging and soon runs into a crash. However, it can properly resolve the rotation of black holes before that. The dynamic behavior of the first-order and second-order scheme are presented in Figure 19, both schemes have the same trajectories until the approaching punctures accumulate errors and start to affect the evolution. Furthermore, Figure 20 shows the evolution of the $L_{2}$ error of the Hamiltonian constraint $H$ and the magnitude of Momentum constraint $|M|$ for both formulations. The first-order formulation can not stabilize the post-merge remnant as it degenerates back to a static single black hole. In contrast, the second-order formulation remains stable and shows a continuous damping of the constraints violation.
To further validate the accuracy of our code, we conducted a benchmark comparison with GRChombo using the same initial physical conditions. Our simulation replicates the same dynamics with the published code, where the black holes complete approximately one and a half orbits before entering the ring-down phase. Minor differences in the puncture positions were observed over time, which may be from the difference in initial lapse conditions—GRChombo uses a conformal-factor-power initial lapse, while we use an averaged initial lapse. Notably, these differences decrease when higher resolutions or higher-order temporal schemes are employed in our simulation, suggesting a qualitative convergence (not shown). A more robust comparison should be made in terms of observables, such as gravitational waveforms. We plan to conduct a more comprehensive study with large-scale benchmarks in future work.
Overall, the rotation and pre-merger phase are accurately captured in our benchmark, and the locations of the punctures are correctly tracked by the (inverse) shift vector guiding the tracers. The simulation of the dynamic behavior is stable, which supports the argument that any constraint violation is “traveling” through the domain, might be reflected at AMR boundaries, but eventually leaves the domain. We avoid the accumulation of numerical errors at specific locations as long as the black holes move.
@@
Figure 20: Evolution of the $L_{2}$ error of the Hamiltonian constraint $H$ and the magnitude of Momentum constraint $|M|$ for the binary black hole scenario for both formulations. The second-order formulation is stable throughout the evolution and shows a continuous damping of the violations after the merge. The first-order formulation can capture the dynamics of black holes correctly and yields only a slight increase in error during the rotation phase. However, it runs into a crash soon after the merge, when the system degenerates back to a single black hole scenario.
@@ 5.4 Context of Experimental Studies
Our application successfully completes a rotating binary black hole merger. It is able to handle both the approach phase as well as the actual ringdown. Our studies focus on the introduction of numerical ingredients within ExaGRyPE. They validate that the software is fully operational and prepared for more challenging benchmarks on larger scales, which will be crucial for further verifying its accuracy and performance. Here, we have to combine all ingredients and notably study a larger computational domain to avoid boundary artifacts (cmp. squeeze-in effect in Figure 19) and obtain realistic measurements. Further stationery tracers (cmp. Section 5.1) then can be used to sample the solution on a huge sphere around the area of interest. With such samples, we can finally calculate $\psi_{4}$ and its mode quantities.
@@ 6 Discussion and Conclusion
To implement this vision and to showcase the applicability and maturity of the design, the ExaGRyPE endeavor had to add further building blocks to the underlying ExaHyPE engine, and we had to run an extensive set of numerical experiments. The experiments demonstrate that our code is fully operational, but also raise further questions, notably around the numerical stability of any composition of ingredients and the philosophical question if first- or second-order formulations are superior. Many aspects of these questions are intrinsically tied to this area of research and have been documented, either implicitly or explicitly, by other codes and groups, too. We consider it a strength of ExaGRyPE that we can quite flexibly study and swap different methodological ingredients to address these fundamental research questions, although the present work does not yet provide answers to many of them. The unique solver coupling features of ExaGRyPE make it an ideal tool for more advanced numerical relativity simulations. For instance, it allows the combination of a finite volume scheme for the matter field with a finite difference scheme for the spacetime background, leading to higher-quality simulations of relativistic hydrodynamics in scenarios such as black hole accretion and neutron star spacetimes. In the same way, undergoing extensions to the particle toolbox—currently used for the tracers—could enable general-relativistic smoothed particle hydrodynamics (SPH) simulations [57, 58, 59].
@@ Appendix B Spacetime Foliation and Evolution Equations
One can already conduct simulations in this stage as $\gamma_{ij},K_{ij},\alpha,\beta^{i}$ with $Z^{a}$ field form a complete evolution system. We further follow [16] and introduce a conformal factor $\phi$ and a conformal metric $\tilde{\gamma}_{ij}$, such that
@@
$\begin{split}&\partial_{t}\hat{\Gamma}^{i}-\beta^{k}\partial_{k}\hat{\Gamma}^{i}-2\alpha\tilde{\gamma}^{ki}\partial_{k}\Theta-\tilde{\gamma}^{kl}\partial_{(k}B_{l)}^{i}-s\frac{1}{3}\tilde{\gamma}^{ik}\partial_{(k}B_{l)}^{l}\\ &\quad\quad-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2\alpha\tilde{\gamma}^{ik}\tilde{\gamma}^{nm}\partial_{k}\tilde{A}_{nm}}+\frac{4}{3}\alpha\tilde{\gamma}^{ij}\partial_{j}K=\frac{2}{3}\tilde{\Gamma}^{i}B_{k}^{k}-\tilde{\Gamma}^{k}B_{k}^{i}\\ &\quad\quad+2\alpha\left(\tilde{\Gamma}_{jk}^{i}\tilde{A}^{jk}-3\tilde{A}^{ij}\frac{P_{j}}{\phi}\right)-2\alpha\tilde{\gamma}^{ki}\left(\Theta\frac{A_{k}}{\alpha}+\frac{2}{3}KZ_{k}\right)\\ &\quad\quad-2\tilde{A}^{ij}A_{j}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}4\alpha\tilde{\gamma}^{ik}D_{k}^{nm}\tilde{A}_{nm}}-2\alpha\kappa_{1}\tilde{\gamma}^{ij}Z_{j}\\ &\quad\quad+2\kappa_{3}\left(\frac{2}{3}\tilde{\gamma}^{ij}Z_{j}B_{k}^{k}-\tilde{\gamma}^{jk}Z_{j}B_{k}^{i}\right),\end{split}$ (58)
@@
$\begin{split}&\partial_{t}A_{k}-\beta^{l}\partial_{l}A_{k}+\alpha^{2}g(\alpha)\left(\partial_{k}K-\partial_{k}K_{0}-2c\partial_{k}\Theta\right)\\ &\qquad+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha g(\alpha)\tilde{\gamma}^{nm}\partial_{k}\tilde{A}_{nm}}=B^{l}_{k}A_{l}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2\alpha g(\alpha)D_{k}^{\ nm}\tilde{A}_{nm}}\\ &\quad\quad\quad\quad\quad\quad-\left[2{\alpha}g(\alpha)+\alpha^{2}g^{\prime}(\alpha)\right]\left(K-K_{0}-2c\Theta\right)A_{k},\end{split}$ (60)
$\begin{split}&\partial_{t}B_{k}^{i}-\beta^{l}\partial_{l}B_{k}^{i}-f\partial_{k}b^{i}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha^{2}\mu\frac{\tilde{\gamma}^{ij}}{\phi}\left(\partial_{k}P_{j}-\partial_{j}P_{k}\right)}\\ &\quad\quad\quad\quad\quad-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha^{2}\mu\tilde{\gamma}^{ij}\tilde{\gamma}^{nl}\left(\partial_{k}D_{ljn}-\partial_{l}D_{kjn}\right)}=B_{k}^{l}B_{l}^{i},\end{split}$ (61)
$\begin{split}&\partial_{t}D_{kij}-\beta^{l}\partial_{l}D_{kij}+s\Big{(}-\frac{1}{2}\tilde{\gamma}_{mi}\partial_{(k}B_{j)}^{m}-\frac{1}{2}\tilde{\gamma}_{mj}\partial_{(k}B_{i)}^{m}\\ &\quad+\frac{1}{3}\tilde{\gamma}_{ij}\partial_{(k}B_{m)}^{m}\Big{)}+\alpha\partial_{k}\tilde{A}_{ij}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha\frac{1}{3}\tilde{\gamma}_{ij}\tilde{\gamma}^{nm}\partial_{k}\tilde{A}_{nm}}\\ &\quad=B_{k}^{l}D_{lij}+B_{j}^{l}D_{kli}+B_{i}^{l}D_{klj}-\frac{2}{3}B_{l}^{l}D_{kij}\\ &\quad-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha\frac{2}{3}\tilde{\gamma}_{ij}D_{k}^{nm}\tilde{A}_{nm}}-A_{k}\left(\tilde{A}_{ij}-\frac{1}{3}\tilde{\gamma}_{ij}\operatorname{tr}\tilde{A}\right),\end{split}$ (62)
$\begin{split}&\partial_{t}P_{k}-\beta^{l}\partial_{l}P_{k}-\frac{1}{3}\phi\left(\alpha\partial_{k}K-\partial_{(k}B_{l)}^{l}\right)-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{1}{3}\alpha\phi\tilde{\gamma}^{nm}\partial_{k}\tilde{A}_{nm}}\\ &\quad=B^{l}_{k}P_{l}+\frac{1}{3}\left(\alpha K-B^{l}_{l}\right)P_{k}+\frac{1}{3}\phi KA_{k}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{2}{3}\alpha\phi D_{k}^{\ nm}\tilde{A}_{nm}}.\end{split}$ (63)
Most of the symbols already appear in the context above and are self-explaining. The $\mathrm{TF}$ index means the Trace-Free part of the quantities, i.e., $R^{\mathrm{TF}}_{ij}=R_{ij}-\frac{1}{3}\gamma_{ij}\gamma^{kl}R_{kl}$. $\operatorname{tr}\tilde{A}=\gamma^{ij}\tilde{A}_{ij}$ is the trace of the conformal traceless extrinsic curvature. There are several red terms added to the evolution systems, using the constraints (50) above for a more symmetric characteristic matrix [15]. Those terms are only employed in the first formulation in ExaGRyPE and are removed in the second-order formulation. The $g(\alpha)$ function is the one defined in the slicing condition in Section 2.1, and the “1+log” slicing corresponds to $g(\alpha)=2/\alpha$. Several new parameters are also introduced for numerical optimization:
@@ Appendix C $\psi_{4}$ in ExaGRyPE
where ${}^{(4)}C_{abcd}$ is the four-dimensional Weyl tensor, and $k^{a}$ and $\bar{m}^{a}$ are two members of a null tetrad. A tetrad is a set of four null vectors that form a complete basis, and we name them $l^{a}$, $k^{a}$, $m^{a}$ and $\bar{m}^{a}$. The tetrad is constructed such that the $l^{a}$ and $k^{a}$ are radial outgoing and ingoing vectors respectively, and only non-vanishing inner products between them are: